A satellite is launched to orbit the Earth at an a
Answer #1

a) orbital period (T):

T^2 =(4*\pi^2/(G*M))*r^3 =(4*\pi^2/(6.67*10^-11*5.97*10^24))* (6.38*10^6)^3 =25746702

=> T=5074.12s=1.4089hr

b)orbital speed v=distance/time =2*\pi*R/T =7900m/s